# captionUnknown document class (or package)

### v1

@@ V Technical Approach
Algorithm 1 Procedure for robot $A\in\mathcal{R}$, a member of robot swarm $\mathcal{R}$, to generate instantaneous relative poses w.r.t. all other agents in swarm $B\in(\mathcal{R}\setminus\{A\})$
$\mathcal{R}$, $A\in\mathcal{R}$, $\{\mathcal{I}_{R}~{}\forall~{}R\in\mathcal{R}\}$, $\ell$
@@
V-A Local Robot Definitions
@@
  • $\mathcal{P}_{R}=\{\mathbf{T}^{R}_{R_{1}},\mathbf{T}^{R}_{R_{2}},\dots,\mathbf{T}^{R}_{R_{N_{R}}}\}$ is an ordered set of $N_{R}$ constant relative poses from $R$’s body frame to $R$’s $k$th antenna’s frame. Additionally, from the information in $\mathbf{T}^{R}_{R_{k}}$, we can succinctly denote the point coordinate of $R$’s $k$th antenna in $R$’s frame as $\mathbf{p}^{R}_{k}$.
Since $W$ is defined with respect to a level ground and gravity vector, the parameters $\langle z^{W}_{R},\alpha^{W}_{R},\beta^{W}_{R}\rangle$ – unlike $\langle x^{W}_{R},y^{W}_{R},\gamma^{W}_{R}\rangle$ – are instantaneously observable and directly measurable from $R$ (e.g., via altimeter or downward facing LiDAR and onboard IMU). Thus, the directly measured values $\tilde{\bm{\kappa}}_{R}=\begin{bmatrix}\tilde{z}^{W}_{R}&\tilde{\alpha}^{W}_{R}&\tilde{\beta}^{W}_{R}\end{bmatrix}^{\top}$ can be continuously monitored by $R$ locally (i.e., without relying on swarm communication). After an initial one-off transmission of $\mathcal{I}_{R}$ to all agents $\mathcal{R}$, agent $R$ only needs to transmit again in the event of (1) wanting to change its commanded altitude/roll/pitch $\hat{\bm{\kappa}}_{R}$, or (2) it locally observes a violation of its previously communicated constraints, specifically:
@@
V-B Inter-Agent Robot Definitions
@@
$d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\triangleq||\mathbf{T}^{A}_{B}\breve{\mathbf{p}}^{B}_{j}-\breve{\mathbf{p}}^{A}_{i}||_{2}$ (2)
where $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\in\mathbb{R}_{\geq 0}$, as distances are non-negative, and $||\cdot||_{2}$ denotes the $L_{2}$ norm. Similarly, let $\tilde{d}^{A_{i}}_{B_{j}}$ be the current (i.e., most recent) noisy measurement between $A$’s $i$th antenna and $B$’s $j$th antenna as measured locally by $A$ [^6] Note that while $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})=d^{{B_{j}}}_{{A_{i}}}(\mathbf{T}^{B}_{A})$ (i.e., the observation model is symmetric), since each agent locally and separately measures antenna distances, measurement noise makes it that generally $\tilde{d}^{A_{i}}_{B_{j}}\neq\tilde{d}^{B_{j}}_{A_{i}}$. For convenience, $\tilde{\mathbf{d}}^{A}_{B}\in\mathbb{R}^{{N_{A}}{N_{B}}}$ denotes the stacked vector of all current pairwise $\tilde{d}^{A_{i}}_{B_{j}}$ measurements.
V-C Noise Model Definition
@@
V-D Optimization Definition
Using our current measurements $\tilde{d}^{A_{i}}_{B_{j}}$, measurement bias correction function $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, and observation model $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, we can formulate the error residual function with respect to antenna pair ${A_{i}}$ and ${B_{j}}$ as:
$e^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\triangleq\underbrace{\Big{(}\tilde{d}^{A_{i}}_{B_{j}}-\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\Big{)}}_{\text{bias adjusted measurement}}-\underbrace{d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})}_{\text{expected measurement}}$ (5)
@@ V-D Optimization Definition
$\min_{\begin{subarray}{c}x^{A}_{B},y^{A}_{B}\in\mathbb{R}\\ \gamma^{A}_{B}\in[-180^{\circ},180^{\circ}]\end{subarray}}f(x^{A}_{B},y^{A}_{B},\gamma^{A}_{B}~{}|~{}\tilde{\mathbf{d}}^{A}_{B},\mathcal{I}_{A},\mathcal{I}_{B},\ell)$ (9)
That is, given our current ranging measurements $\tilde{\mathbf{d}}^{A}_{B}$, our known information tuples $\mathcal{I}_{A}$ and $\mathcal{I}_{B}$, and some loss function $\ell$, our 3D pose can be found as 3-DoF minimization of our free variables $\langle x^{A}_{B},y^{A}_{B},\gamma^{A}_{B}\rangle$.
Using Equation 9, Algorithm 1 outlines a procedure independently followed by each agent $A$ in a swarm $\mathcal{R}$ to individually estimate $\mathbf{T}^{A}_{B}$ between itself and all other agents $B\in(\mathcal{R}\setminus\{A\})$. Observe that one-off communication (i.e., transmit) only occurs on Line 4, when an agent locally detects it has violated its a priori constraints – this is then used by other agents on Line 7 (i.e., invalid) to skip calculating $\mathbf{T}^{A}_{B}$ for agents with violated constraints.
TABLE II: Evaluation of positional error in meters with different $z$-constraints. In practice, the second row (i.e., a priori constraints) will be used, but the small difference between z_comm, z_meas, and z_true shows these constraints are not a major source of error. See Section VI-A.
Abs Position Error [m] Scenarios (Section VI-A)
@@
$\blacksquare$
z_comm 0.26 0.45 0.09 0.28 0.87 0.14 0.25 0.64 0.13
$\blacksquare$
z_meas 0.26 0.44 0.09 0.27 0.87 0.14 0.24 0.66 0.14
$\blacksquare$
z_true 0.25 0.44 0.10 0.26 0.87 0.15 0.22 0.64 0.14
(a) Shows Table II’s Trial 2 alongside the commanded, measured, and ground truth altitude, where the colors correspond with Table II. While the TeraRanger Evo does not perfectly track ground truth, it stays within the expected commanded tolerance providing a reasonable local measurement of altitude.
(b) Top plot shows calibrated ranging error (i.e., mean bias subtracted) w.r.t. relative yaw for the 1m above and level flights (i.e., Trial 4 and 5 respectively). Observe the error of 1m above flight remains consistent, while the level flight has anomalous error peaks corresponding to specific relative yaws. The lower left and right diagrams show the relative pose of the agents at the two most prominent peaks, where the path between antennas $A_{1}$ and $B_{5}$ is eclipsed by antennas $B_{3}$ and $B_{1}$ respectively. While this NLOS effect is consistent with our previous 2D work [18], we note the more complex 3D relationship resolves the occlusion with varying elevation.
@@
TABLE III: Evaluation of positional error in meters between algorithms. The final row represents the proposed approach. See Section VI-B.
TABLE IV: Evaluation of yaw error in meters between algorithms. The final row represents the proposed approach. See Section VI-B.
Abs Position Error [m] Scenarios (Section VI-B)
@@
$\blacksquare$
1.14 3.29 1.12 2.15 3.29 1.03 1.24 3.60 1.01 2.30 3.57 1.02 1.30 3.24 1.16
$\blacksquare$
0.44 1.64 0.30 0.95 2.58 0.90 0.85 3.10 0.75 0.85 2.69 0.82 1.44 2.88 0.90
$\blacksquare$
0.29 0.91 0.14 0.27 1.02 0.14 0.32 1.03 0.20 0.28 0.78 0.14 0.29 0.71 0.13
$\blacksquare$
0.22 0.90 0.17 0.21 0.98 0.16 0.30 0.97 0.21 0.23 0.75 0.15 0.22 0.63 0.14
Abs Heading Error [deg] Scenarios (Section VI-B)
@@
$\blacksquare$
8.0 47.5 5.5 7.7 177.7 12.8 8.5 57.5 6.1 10.1 111.4 10.7 10.4 47.5 7.6
$\blacksquare$
7.8 47.7 5.7 7.8 177.3 12.8 8.5 57.3 6.1 10.3 111.2 11.0 10.4 46.5 7.6
$\blacksquare$
7.6 42.4 5.1 7.6 172.6 12.5 8.3 58.4 6.1 10.4 122.0 11.2 10.6 48.5 7.6
$\blacksquare$
8.6 46.7 5.8 8.4 172.8 12.8 8.7 59.8 6.5 10.7 111.8 11.2 11.0 42.3 7.7
TABLE IV: Evaluation of yaw error in meters between algorithms. The final row represents the proposed approach. See Section VI-B.
(a) Proposed approach (red) generally outperforms other approaches with respect to absolute position error – only a few select points being the exception.
(b) Proposed approach (red) generally equivalent to other approaches with respect to absolute heading error, but is outperformed at a few select points.
@@
Figure 5: Error vs Time plots for Trial 2 in Table IV and IV respectively. The line color corresponds with the algorithm’s color column in the associated tables. Each colored dashed line represent that algorithm’s overall mean in that trial.
VI Experimental Results
@@
We present two sets of experiments: UAV (Section VI-A) and UGV (Section VI-B). We first verify the 3D noise properties, flight hardware, and show altitude/roll/pitch constraints can be sufficiently locally monitored on a UAV. Once this is verified, we confidently attach surrogate UAVs at different altitudes to UGVs – this allows us to collect data more efficiently to evaluate our approach.
VI-A UAV Experiments
@@
Hardware: Two agents, $A$ and $B$, were used. $A$, a surrogate quadcopter, was a stationary Turtlebot2 UGV with 4 UWB sensors evenly distributed in the $xy$-plane 0.37m from the center (Same robot used in prior work [18]). As pictured in Figure 1, $B$ is a custom hexrotor equipped with 6 UWB sensors evenly distributed in the $xy$-plane 0.31m from the center. Additionally, $B$ is equipped with a downward facing TeraRanger Evo 15m LiDAR for altitude measurements [33].
@@
Trials: Five experimental trials were conducted in a large 10m×$\times$10m mocap space. For all trials, $A$ was placed on a table for additional height and kept stationary, while $B$ flew around. In the first three trials, $B$ took off and flew simple line patterns at different altitudes (see Table II and Figure 3(a)). For the latter two trials, $B$ flew tight circles in the center of the room – first 1m above and then level with $A$ (see Figure 3(b)). Each trial had a max speed of 1m/s and lasted approximately 3-5min (i.e., the length of a battery charge).
@@
Discussion of Results: Results are shown in Figure 4 and Table II. The main takeaways are: (1) Verified UWB measurements were not corrupted by EMI of nearby propeller motors. (2) Confirmed the robot’s body and other antennas cause significant NLOS measurement noise and bias when eclipsing a measurement path. In our previous 2D work [18], obstruction was only a function of relative azimuth and yaw and was addressed by a custom weighting function. In 3D, the addition of $\langle z,\alpha,\beta\rangle$ makes this relationship overall more complex – that being said, when agents are not in the same $z$-plane and remain level, these NLOS effects go away (see Figure 3(b)). Future work will explore modeling this entire relationship as a neural net directly within $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, but for this work a robust Huber loss function (see Section VI-B) is sufficient for agents when $\Delta z\neq 0$. (3) Confirmed the onboard downward facing LiDAR and IMU can locally monitor altitude/roll/pitch accurately with only small variation from the mocap ground truth. Figure 3(a) shows Table II’s Trial 2, convincing us the TeraRanger Evo provides a reasonable local measurement of altitude. Table II shows the results of the proposed algorithm when $z$ is free (z_free), or fixed to the commanded (z_comm), LiDAR measured (z_meas), or true (z_true) value. The small APE difference between the latter three shows variation within our a priori tolerances is not a major source of error, and we can detect if we leave said bounds. This validates our minimal communication model. (4) Together, these takeaways give us confidence that surrogate UAVs attached at different altitudes to UGVs are sufficient to fairly evaluate our approach in Section VI-B.
VI-B UGV Experiments

### v3

@@ IV Technical Approach
Algorithm 1 Procedure for robot $A\in\mathcal{R}$, a member of robot swarm $\mathcal{R}$, to generate instantaneous relative poses w.r.t. all other agents in swarm $B\in(\mathcal{R}\setminus\{A\})$. $A$ communicates violated assumptions with $\texttt{transmit}(A)$. $A$ locally tracks which agents have communicated violated assumptions, and checks $B$’s last received status with $\texttt{invalid}(B)$.
$\mathcal{R}$, $A\in\mathcal{R}$, $\{\mathcal{I}_{R}~{}\forall~{}R\in\mathcal{R}\}$, $\ell$, $\bar{d}^{A_{i}}_{B_{j}}$
@@
IV-A Pose Parameterization
Given a set of 3D reference frames $\mathcal{F}$, consider any pair of frames $A\in\mathcal{F}$ and $B\in\mathcal{F}$. Let $\mathbf{T}^{A}_{B}\in SE(3)$ be the relative pose from frame $A$ to frame $B$. We note $\mathbf{T}^{A}_{B}$ is a 6-DoF value that can be equivalently parameterized as the 2-tuple $\langle\mathbf{R}^{A}_{B},\mathbf{t}^{A}_{B}\rangle$ or 6-tuple $\langle x^{A}_{B},y^{A}_{B},z^{A}_{B},\alpha^{A}_{B},\beta^{A}_{B},\gamma^{A}_{B}\rangle$, where $\mathbf{R}^{A}_{B}\in SO(3)$ is the relative rotation between $A$ and $B$ parameterized by relative roll $\alpha^{A}_{B}$, pitch $\beta^{A}_{B}$, and yaw $\gamma^{A}_{B}$ and $\mathbf{t}^{A}_{B}\in\mathbb{R}^{3}$ is a relative translation vector between $A$ and $B$ parameterized by relative $x^{A}_{B}$, $y^{A}_{B}$, and $z^{A}_{B}$.
IV-B Local Robot Definitions
@@
  • $\mathcal{P}_{R}=\{\mathbf{T}^{R}_{R_{1}},\mathbf{T}^{R}_{R_{2}},\dots,\mathbf{T}^{R}_{R_{N_{R}}}\}$ is an ordered set of $N_{R}$ constant relative poses from $R$’s body frame to $R$’s $k$th antenna’s frame. From $\mathbf{T}^{R}_{R_{k}}$, we can succinctly denote the location of $R$’s $k$th antenna in $R$’s frame as point $\mathbf{p}^{R}_{k}$, where $\breve{\mathbf{p}}^{R}_{k}$ is the equivalent homogeneous point.
Since $W$ is defined with respect to a level ground and gravity vector, the parameters $\langle z^{W}_{R},\alpha^{W}_{R},\beta^{W}_{R}\rangle$ – unlike $\langle x^{W}_{R},y^{W}_{R},\gamma^{W}_{R}\rangle$ – are instantaneously observable and directly measurable from $R$ (e.g., via altimeter or downward facing LiDAR and onboard IMU). Thus, the directly measured values $\tilde{\bm{\kappa}}_{R}=\begin{bmatrix}\tilde{z}^{W}_{R}&\tilde{\alpha}^{W}_{R}&\tilde{\beta}^{W}_{R}\end{bmatrix}^{\top}$ can be continuously monitored by $R$ locally (i.e., without relying on swarm communication). After an initial transmission of $\mathcal{I}_{R}$ to all agents $\mathcal{R}$, agent $R$ only needs to transmit again in the event of (1) wanting to change its commanded altitude/roll/pitch $\hat{\bm{\kappa}}_{R}$, or (2) it locally observes a violation of its previously communicated constraints, specifically:
@@
IV-C Inter-Agent Robot Definitions
@@
$d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\triangleq\lVert\mathbf{T}^{A}_{B}\breve{\mathbf{p}}^{B}_{j}-\breve{\mathbf{p}}^{A}_{i}\rVert_{2}$ (2)
where $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\in\mathbb{R}_{\geq 0}$, as distances are non-negative, and $\lVert\cdot\rVert_{2}$ denotes the $L_{2}$ norm. Similarly, let $\tilde{d}^{A_{i}}_{B_{j}}$ be the current (i.e., most recent) noisy measurement between $A$’s $i$th antenna and $B$’s $j$th antenna as measured locally by $A$ [^8] Note that while $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})=d^{{B_{j}}}_{{A_{i}}}(\mathbf{T}^{B}_{A})$ (i.e., the observation model is symmetric), since each agent locally and separately measures antenna distances, measurement noise makes it that generally $\tilde{d}^{A_{i}}_{B_{j}}\neq\tilde{d}^{B_{j}}_{A_{i}}$. For convenience, $\tilde{\mathbf{d}}^{A}_{B}\in\mathbb{R}^{{N_{A}}{N_{B}}}$ denotes the stacked vector of all current pairwise $\tilde{d}^{A_{i}}_{B_{j}}$ measurements.
IV-D Noise Model Definition
@@
IV-E Optimization Definition
Consider the following error residual function between the measured and expected distances for antenna pair ${A_{i}}$ and ${B_{j}}$ with respect to some $\mathbf{T}^{A}_{B}$, a relative pose between $A$ and $B$ (and our eventual decision variable):
$e^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\triangleq\underbrace{\Big{(}\tilde{d}^{A_{i}}_{B_{j}}-\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})\Big{)}}_{\text{bias adjusted measurement}}-\underbrace{d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})}_{\text{observation model}}$ (5)
where $\tilde{d}^{A_{i}}_{B_{j}}$ is the current measurements, $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$ is a measurement bias correction function (current implementation in Section V-B), and $d^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$ the known observation model.
@@ IV-E Optimization Definition
$\min_{\begin{subarray}{c}x^{A}_{B},y^{A}_{B}\in\mathbb{R}\\ \gamma^{A}_{B}\in[-180^{\circ},180^{\circ}]\end{subarray}}f(x^{A}_{B},y^{A}_{B},\gamma^{A}_{B}~{}|~{}\tilde{\mathbf{d}}^{A}_{B},\mathcal{I}_{A},\mathcal{I}_{B},\ell,\bar{d}^{A_{i}}_{B_{j}})$ (9)
Using Equation 9, Algorithm 1 outlines a procedure independently followed by each agent $A$ in a swarm $\mathcal{R}$ to individually estimate $\mathbf{T}^{A}_{B}$ between itself and all other agents $B\in(\mathcal{R}\setminus\{A\})$. Observe that event-based communication (i.e., transmit) only occurs on Line 4, when an agent locally detects it has violated its a priori constraints – other agents use this information on Line 7 (i.e., invalid) to skip calculating $\mathbf{T}^{A}_{B}$ for agents with violated constraints.
TABLE III: Evaluation of positional error in meters with different $z$-constraints. In practice, the second row (i.e., a priori constraints) will be used, but the small difference between z_comm, z_meas, and z_true shows these constraints are not a major source of error. See Section V-A.
Abs Position Error [m] Scenarios (Section V-A)
@@
$\blacksquare$ z_comm 0.26 0.45 0.09 0.28 0.87 0.14 0.25 0.64 0.13
$\blacksquare$ z_meas 0.26 0.44 0.09 0.27 0.87 0.14 0.24 0.66 0.14
$\blacksquare$ z_true 0.25 0.44 0.10 0.26 0.87 0.15 0.22 0.64 0.14
(a) Shows Table III’s Trial 2 alongside the commanded, measured, and ground truth altitude, where the colors correspond with Table III. While the TeraRanger Evo does not perfectly track ground truth, it stays within the expected commanded tolerance of ±$\pm 0.10$m, providing a reasonable local measurement of altitude.
(b) Top plot shows calibrated ranging error (i.e., mean bias subtracted) with respect to relative yaw for the 1m Above and Level flights (i.e., Trial 4 and 5 respectively). Observe the error of 1m Above flight remains consistent, while the Level flight has anomalous error peaks, mostly within the ±$\pm 0.15$m range, corresponding to specific relative yaws. The lower left and right diagrams show the relative pose of the agents at the two most prominent peaks (i.e., with error greater than $0.45$m), where the path between antennas $A_{1}$ and $B_{5}$ is eclipsed by antennas $B_{3}$ and $B_{1}$ respectively. While this NLOS effect is consistent with our previous 2D work [10], we note the more complex 3D relationship resolves the occlusion with varying elevation.
@@
TABLE IV: Evaluation of positional error in meters between algorithms. The final row (red) represents the proposed approach. See Section V-B.
Abs Position Error [m] Experimental Scenarios (Section V-B)
@@
$\blacksquare$ 1.14 3.29 1.12 2.15 3.29 1.03 1.24 3.60 1.01 2.30 3.57 1.02 1.30 3.24 1.16
$\blacksquare$ 0.44 1.64 0.30 0.95 2.58 0.90 0.85 3.10 0.75 0.85 2.69 0.82 1.44 2.88 0.90
$\blacksquare$ 0.29 0.91 0.14 0.27 1.02 0.14 0.32 1.03 0.20 0.28 0.78 0.14 0.29 0.71 0.13
$\blacksquare$ 0.22 0.90 0.17 0.21 0.98 0.16 0.30 0.97 0.21 0.23 0.75 0.15 0.22 0.63 0.14
TABLE V: Evaluation of yaw error in degrees between algorithms. The final row (red) represents the proposed approach. See Section V-B.
Abs Heading Error [deg] Experimental Scenarios (Section V-B)
@@
$\blacksquare$ 8.0 47.5 5.5 7.7 177.7 12.8 8.5 57.5 6.1 10.1 111.4 10.7 10.4 47.5 7.6
$\blacksquare$ 7.8 47.7 5.7 7.8 177.3 12.8 8.5 57.3 6.1 10.3 111.2 11.0 10.4 46.5 7.6
$\blacksquare$ 7.6 42.4 5.1 7.6 172.6 12.5 8.3 58.4 6.1 10.4 122.0 11.2 10.6 48.5 7.6
$\blacksquare$ 8.6 46.7 5.8 8.4 172.8 12.8 8.7 59.8 6.5 10.7 111.8 11.2 11.0 42.3 7.7
@@
Figure 5: Error vs Time plots for Trial 2 in Table V and V respectively. The line color corresponds with the algorithm’s color column in the associated tables. Each colored dashed line represent that algorithm’s overall mean in that trial.
Figure 6: Shows the corresponding absolute position error for Trial 1 when each approach (colors correspond with Table V) is run with the incorrect altitude $z$-constraint $\hat{z}^{A}_{B}\triangleq-1.25\text{m}+\Delta z$, where $-1.25$m is the true relative altitude between $A$ and $B$ (Figure 3(a)) and $\Delta z$ is relative altitude error (i.e., the $x$-axis value). The green/blue lines are level since $z$ is unconstrained in these methods, and thus not depended on the $x$-axis. Although purple also has $z$ unconstrained, without Huber loss, it is hyper-sensitive to altitude initialization (i.e., $x$-axis value). Comparing our proposed (red) to the approach without el_bias (orange), we see red outperforms orange even when operating with maximal altitude error within the reasonable error tolerances (i.e., ±$\pm 0.10$m) demonstrated in Figure 4(a).
V Experimental Results
@@
We present two sets of experiments: UAV (Section V-A) and UGV (Section V-B). We first verify the 3D noise properties, flight hardware, and show altitude/roll/pitch constraints can be sufficiently locally monitored on a UAV. Once this is verified, we attach surrogate UAVs at different altitudes to UGVs – this allows us to efficiently collect large amounts of data to evaluate our approach.
V-A UAV Experiments
@@
Hardware: Two agents, $A$ and $B$, were used. $A$, a surrogate quadcopter, was a stationary Turtlebot2 UGV with 4 UWB sensors evenly distributed in the $xy$-plane 0.37m from the center (i.e., same robot in prior work [10]). $B$ is a custom hexrotor UAV equipped with 6 UWB sensors evenly distributed in the $xy$-plane 0.31m from the center (Figure 1). Additionally, $B$ is equipped with a downward facing TeraRanger Evo 15m LiDAR for altitude measurements [24].
@@
Trials: Five experimental trials were conducted in a large 10m×$\times$10m mocap space. For all trials, $A$ was placed on a table for additional height and kept stationary, while $B$ flew around. In the first three trials, $B$ took off and flew simple line patterns at different altitudes (see Table III and Figure 4(a)). For the latter two trials, $B$ flew tight circles in the center of the room – first 1m Above and then Level with $A$ (see Figure 4(b)). Each trial had a max speed of 1m/s and lasted approximately 3-5min (i.e., the length of a battery charge).
@@
Discussion of Results: Results are shown in Table III and Figure 4. The key takeaways are as follows: (1) Demonstrated UWB measurements were not corrupted by proximity to propeller motors (e.g., electromagnetic interference). This gives us confidence measurements in Section V-B are comparable. (2) Showed the 2D NLOS concerns of prior work [10] are mitigated when level 3D agents are not in the same $z$-plane (Figure 4(b)). (3) Confirmed the onboard downward facing LiDAR and IMU can locally monitor altitude/roll/pitch accurately with only small variation from the mocap ground truth. Figure 4(a) shows data from Table III’s Trial 2, demonstrating the TeraRanger Evo provides a reasonable local measurement of altitude. (4) Confirmed varying within our a priori tolerances are not a major source of APE error, supported by the small APE differences between the latter three rows of Table III. Specifically, this table shows the results of the proposed algorithm when $z$ is free (z_free), or fixed to the commanded (z_comm), LiDAR measured (z_meas), or true (z_true) value. Together, takeaways (3) and (4) show that hardware realistically allows for an altitude tolerance of ±$\pm 0.10$m, meaning under realistic flight conditions, our minimal communication model introduces only a small amount of additional error (i.e., compare z_comm and z_true in Table III). Furthermore, all the takeaways together indicate our extensive UGV tests are sufficient to fairly evaluate our approach in Section V-B.
V-B UGV Experiments